# Synthetic bootstrapped pretraining

### v2

@@ A.2 Ablation on data mixture ratio
When performing joint training on a mixture of real and synthesized documents for the final SBP run, a natural question arises: how much fraction of synthesized documents to include. In §4, we discussed that we utilized $\|{{\mathcal{S}}_{\text{pretrain}}}\|=$75B for the 200B-scale experiment and $\|{{\mathcal{S}}_{\text{pretrain}}}\|=$125B for the 1T-scale experiment. In this section, we present ablation experiments for this design choice.
@@
Figure 7: SBP performance with varying synthetic tokens at 1T-scale.
@@ Discussion
From this analysis, we can observe a general pattern: the best-performing model is achieved when pretraining is conducted on a mixture of real and synthetic data. Real internet data has higher quality and therefore merits more repetition. However, as repetition yields diminishing returns, synthetic data could offer another source of signal that real data cannot capture. In contrast, distillation-based research typically finds that training purely on synthetic data yields significantly higher training efficiency. However, this finding is obscured by the fact that such a model eventually converges to the capability of the teacher LM. This contrast reveals that the SBP mechanism does not generate a compressed and denoised representation of knowledge that is more efficient for LM to learn. Instead, it offers an additional source of improvement that real data alone cannot capture.
@@ B.1 Analyze concepts in documents
The distribution of concept domains and types in Table 4 and 5 underscores the multidimensional nature of the knowledge space under consideration. The domains encompass macro-level sociocultural phenomena, such as Culture, where topics range from inter-community conflict in Nigeria to immigration policy and interracial dating and bias, alongside micro-level issues of individual health and wellbeing, as exemplified in Health. In parallel, the typological classification reveals not only subject matter but also modes of conceptual engagement: Methods comprise formalized procedures (multidimensional poverty measurement, commercial real estate appraisal), Events capture historically situated crises (Mediterranean migrant crisis, BP oil spill nationalization), and Comparisons and Analyses facilitate interpretive framing through juxtapositions (cancer suffering: individual vs. family) and evaluative inquiries (Manchester United player analysis). Collectively, this taxonomy illustrates not only topical diversity but also a spectrum of cognitive orientations.
@@ C.1 Two epochs validation
When designing the oracle experiment for 1T-scale, we noted that we use 482B tokens repeated twice as a proxy for training on 1T unique tokens. This is because the DCLM-baseline (Li et al., 2024) dataset contains 80% duplicates, which hinders our evaluation. We validate our choice by scaling down the experiment to a 400B scale, where we had sufficiently many unique tokens. As seen in Table 10, 200B tokens repeated twice yield nearly identical performance to 400B unique tokens. This finding is consistent with the observation from Muennighoff et al. (2023) where repetition up to 4 times yields nearly no performance degradation.
@@ C.2 Model Scaling
Table 12: 200B-scale experiments with model scaling. The first three columns are identical to Table1. The last column shows the performance of training a 6B model under 200B training token budget with 10B unique tokens.

### v3

@@ A.2 Ablation on data mixture ratio
When performing joint training on a mixture of real and synthesized documents for the final SBP run, a natural question arises: how much fraction of synthesized documents to include. In §4, we discussed that we utilized $\|{{\mathcal{S}}_{\text{pretrain}}}\|=$75B for the 200B-scale experiment, $\|{{\mathcal{S}}_{\text{pretrain}}}\|=$125B for the 1T-scale (3B) experiment, and $\|{{\mathcal{S}}_{\text{pretrain}}}\|=$250B for the 1T-scale (6B) experiment. In this section, we present ablation experiments for this design choice.
@@
Figure 7: SBP performance with varying synthetic tokens at 1T-scale (3B).
1T-scale (6B)
We also performed the mixture ratio sweep for the 6B model at the 1T-scale, evaluating $\|{{\mathcal{S}}_{\text{pretrain}}}\|\in\{$0B, 125B, 250B}\}. As shown in Figure 8, we find that the optimal amount of synthetic data is around 250B, which is higher than the optimal 125B observed for the 3B model.
Figure 8: SBP performance with varying synthetic tokens for the 6B model at 1T-scale.
@@ Discussion
From this analysis, we can observe a general pattern: the best-performing model is achieved when pretraining is conducted on a mixture of real and synthetic data. Furthermore, the optimal ratio of synthetic data appears to increase with model size (from  12.5% for 3B to  25% for 6B). Real internet data has higher quality and therefore merits more repetition. However, as repetition yields diminishing returns, synthetic data could offer another source of signal that real data cannot capture. In contrast, distillation-based research typically finds that training purely on synthetic data yields significantly higher training efficiency. However, this finding is obscured by the fact that such a model eventually converges to the capability of the teacher LM. This contrast reveals that the SBP mechanism does not generate a compressed and denoised representation of knowledge that is more efficient for LM to learn. Instead, it offers an additional source of improvement that real data alone cannot capture.
A.3 Random pairs and embedding analysis
To verify that SBP relies on learning specific inter-document correlations rather than generic data augmentation, we conduct an ablation study where we train the synthesizer on random document pairs. We measure the semantic similarity between the seed document $d_{1}$} and the target document (either $d_{2}$} or the synthesized output) using the Qwen3-Embedding-0.6B model.
Table 4: Embedding similarity statistics. “Paired documents” refers to the SBP training pairs found by nearest neighbor search. “Random documents” refers to randomly paired documents. “Generated documents (SBP)” refers to the synthetic data generated by the SBP model at 200B-scale (3B). “Generated documents (Random)” refers to the synthetic data generated by the model trained on random pairs. All comparisons are based on the 10B dataset.
Statistic Paired docs Random docs Generated (SBP) Generated (Random)
Mean 0.79 0.15 0.66 0.32
As shown in Table 4, the paired documents used in SBP exhibit high semantic similarity (0.79), whereas random documents have minimal similarity (0.15). Crucially, the documents generated by the SBP synthesizer maintain a high degree of relevance (0.66) to the seed document. In contrast, the model trained on random pairs produces outputs with significantly lower relevance (0.32). This result confirms that the SBP synthesizer successfully learns to preserve semantic relevance from the training pairs, a property that is absent when training on random associations.
@@ B.1 Analyze concepts in documents
The distribution of concept domains and types in Table 5 and 6 underscores the multidimensional nature of the knowledge space under consideration. The domains encompass macro-level sociocultural phenomena, such as Culture, where topics range from inter-community conflict in Nigeria to immigration policy and interracial dating and bias, alongside micro-level issues of individual health and wellbeing, as exemplified in Health. In parallel, the typological classification reveals not only subject matter but also modes of conceptual engagement: Methods comprise formalized procedures (multidimensional poverty measurement, commercial real estate appraisal), Events capture historically situated crises (Mediterranean migrant crisis, BP oil spill nationalization), and Comparisons and Analyses facilitate interpretive framing through juxtapositions (cancer suffering: individual vs. family) and evaluative inquiries (Manchester United player analysis). Collectively, this taxonomy illustrates not only topical diversity but also a spectrum of cognitive orientations.
@@ C.1 Two epochs validation
When designing the oracle experiment for 1T-scale, we noted that we use 482B tokens repeated twice as a proxy for training on 1T unique tokens. This is because the DCLM-baseline (li2024datacomplm) dataset contains 80% duplicates, which hinders our evaluation. We validate our choice by scaling down the experiment to a 400B scale, where we had sufficiently many unique tokens. As seen in Table 11, 200B tokens repeated twice yield nearly identical performance to 400B unique tokens. This finding is consistent with the observation from muennighoff2023scaling where repetition up to 4 times yields nearly no performance degradation.
@@ C.2 Model Scaling
Table 13: 200B-scale experiments with model scaling. The first three columns are identical to Table1. The last column shows the performance of training a 6B model under a 200B training token budget with 10B unique tokens.